OpenStudy (anonymous):

Adrian's CD player can hold six disks at a time and shuffles all of the albums and their songs. If he has thirteen CD's, how many different combinations of six CD's can he put in the player?

OpenStudy (kropot72):

The number of combinations of n different things taken r at a time is: \[\large C(n, r)=\frac{n!}{r!(n-r)!}\] In this case n = 13 and r = 6. Just plug these into the formula.

OpenStudy (anonymous):

I do not know how to work the problem out.

OpenStudy (kropot72):

\[\large C(13, 6)=\frac{13\times12\times11\times10\times9\times8}{6\times5\times4\times3\times2\times1}=you\ can\ calculate\]